The interior of a refrigerator having inside dimensions of 0.5 m × 0.5 m

base area and 1 m height, is to be maintained at 6°C. The walls of the refrigerator are constructed

of two mild steel sheets 3 mm thick (k = 46.5 W/m°C) with 50 mm of glass wool insulation (k =

0.046 W/m°C) between them. If the average heat transfer coefficients at the inner and outer

surfaces are 11.6 W/m2°C and 14.5 W/m2°C respectively, calculate :

(i) The rate at which heat must be removed from the interior to maintain the specified

temperature in the kitchen at 25°C, and

(ii) The temperature on the outer surface of the metal sheet.


1
Expert's answer
2021-09-07T09:51:17-0400

The variables that we have are:


Thickness of each mild steel refrigerator wall, La = Lc = 3 mm = 0.003 m

Thickness of glass wool insulation wall, Lb = 50 mm = 0.050 m

Temperature of hot fluid (outside), "t_{o} = 25\u00b0C"

Temperature of cold fluid (inside of refrigerator), "t_{ref} = 6\u00b0C"

Thermal conductivity of mild steel, "k_a=k_c= 46.5 \cfrac{W}{m°C}"

Thermal conductivity of glass wool insulation, "k_b= 0.046 \cfrac{W}{m°C}"

Heat transfer coefficients: Hot fluid (outside), "h_{o} = 11.6 \cfrac{W}{m^2°C}"

Cold fluid (inside of refrigerator), "h_{ref} = 14.5\cfrac{W}{m^2°C}"


First, we use the relation to find the rate at which heat must be removed from the interior to maintain the specified temperature in the kitchen at 25°C as:


"q = UA(t_{o} – t_{i})= \cfrac{A(t_{o} – t_{i})}{\cfrac{1}{h_{o} }+\cfrac{L_a}{k_a}+\cfrac{L_b}{k_b}+\cfrac{L_c}{k_c}+\cfrac{1}{h_{i} }}" 


Before we proceed, we have to find the area of the refrigerator as


 "A=2\times(0.5\,m)(0.5\,m)+4\times(1\,m)(0.5\,m)=2.5\,m^2"


We substitute in the formula for q and find:


"q = \cfrac{(2.5\,m^2)(25 – 6)°C}{\cfrac{1}{11.6 \frac{W}{m^2°C} }+\cfrac{0.003\,m}{46.5 \frac{W}{m°C}}+\cfrac{0.050\,m}{0.046 \frac{W}{m°C}}+\cfrac{0.003\,m}{46.5 \frac{W}{m°C}}+\cfrac{1}{ 14.5\frac{W}{m^2°C} }}"


"\\ \implies q= 38.24 W"


After that, we can also establish another equation for the heat loss on the outside to find the temperature of the shell:

"q = h_{o}A (t_{o} – t_{1}) \implies t_{1} =t_{o} - \cfrac{q}{h_{o}A}"


We substitute and we can confirm the temperature of the outside surface of the metal sheet:


"t_{1}=25°C-\cfrac{38.24\,W}{(11.6 \frac{W}{m^2°C})(2.5\,m^2)} =23.68 °C"


In conclusion, (i) the rate of heat loss of the refrigerator at which heat must be removed from the interior to maintain the specified temperature in the kitchen at 25°C is 38.24 W, and (ii) the temperature of the outer surface of the metal sheet is 23.68 °C.


Reference:

  • Sears, F. W., & Zemansky, M. W. (1973). University physics.